OpenStudy (itrymath):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = x - 3. (3 points) y = 3x - 14 y = x + 2 y = -3x + 22 y = - x + 8

OpenStudy (itrymath):

Choose the slope-intercept equation of the line that passes through the point (6, 4) and is parallel to y = x - 3. (3 points) y = 3x - 14 y = 1/3x + 2 y = -3x + 22 y = -1/3 x + 8

OpenStudy (itrymath):

@Will.H

OpenStudy (itrymath):

@steve816 @563blackghost @TGstudios @mathmate

OpenStudy (itrymath):

@TGstudios hey whats up

OpenStudy (tgstudios):

not sure, very confusing, but the formula in the question has a slope of 1/1 and start at (0,3), which means a line parellel to it would have to have a slope of 1 also...

OpenStudy (tgstudios):

I'm not sure which answer, hope I helped just a little bit... :/

OpenStudy (itrymath):

@Will.H

OpenStudy (will.h):

Parallel to means they have same slope. Therefore the slope is 1 And passes through point (6,4) Y - Y1 = m(x-x1) Y - 4 = 1(x - 6) Y - 4 = x -6 Y = x -2

OpenStudy (itrymath):

@Will.H um thats not an answer choice

OpenStudy (will.h):

I noticed. You should report that question as a fraud. And they tended to want the right answer as B

OpenStudy (itrymath):

what do you mean there has to be a right answer

OpenStudy (will.h):

I solved it correctly. I guarantee. It's just that the question indicates error. Report that to ur instructor and they'll give u appreciation

OpenStudy (itrymath):

but this is a test

OpenStudy (itrymath):

lol

OpenStudy (will.h):

Because it looks similar to the one we found. They seemed to have the slope as -1 instead the real one is is 1 that was their mistake. So that's why they appeared to have choice B. Which seems like the right one based on their crap

OpenStudy (will.h):

Again. Put ur choice as B. And report that question. Show them my work as yours. And you'll be rewarded lol

OpenStudy (itrymath):

@Will.H okay, but 1 question

OpenStudy (will.h):

Shoot